Seti: Kurze Wartungsarbeiten
Diskutiere Seti: Kurze Wartungsarbeiten im TecCentral Power Cruncher Team Forum im Bereich Informationen & Feedback; Außerdem war das mit den Serverausfällen schonmal schlimmer... Ab Herbst letzen Jahres lief es doch alles gut, ein paar geplante Ausfälle gab es ja wegen ...
- 12.04.2006, 12:31 Seti: Kurze Wartungsarbeiten #9
Jetzt hast Du den Idefix aber beleidigt
- Registrieren bzw. einloggen, um diese und auch andere Anzeigen zu deaktivieren
- 17.05.2006, 10:06 Seti: Kurze Wartungsarbeiten #10
Aufgrund eines Problemes waren gestern Abend keine Up-/Downloads möglich...
May 16, 2006
A cascade of problems today caused inability to return results and later, the inability to download new work. One of the data servers lost the ability to see the upload directory. After reboot, it temporarily lost the ability to see the entire network.
- 30.05.2006, 17:25 Seti: Kurze Wartungsarbeiten #11
Bei Seti erfährt man immerhin:
May 30, 2006
The projects have been temporarily stopped in order to diagnose problems that have arisen with file upload handling. This outage will continue into our regularly scheduled maintenance outage today. We expect the projects to be restarted by 21:00 UTC. Data servers and web pages are offline until our maintenance is completed. Thank you for your patience.
- 30.05.2006, 22:11 Seti: Kurze Wartungsarbeiten #12
Jetzt läuft Seti aber... Habe eben meine fertigen WUs hochgeladen bekommen, kam auch erst von der Spätschicht.
May 30, 2006
At 13:30 UTC yesterday the SETI@home upload/download server failed. We've solved the problem, so you should be able to upload results and download new work at this time. There will probably be a long period where the servers are overloaded with backlogged requests. Thank you for your patience.
Bei Einstein ist immer noch alles dunkel, aber immerhin kann die Kiste wieder durchlaufen...
- 08.09.2006, 23:23 Seti: Kurze Wartungsarbeiten #13
Bei Seti@home gibt es im Moment (und vielleicht fürs ganze Wochende) keine WUs aufgrund von Wartungsarbeiten an der Datenbank (neuer Server).
September 8, 2006 - 22:00 UTC
We are out of work to send out, and probably will be for the whole weekend. Here's why:
Recently Sun donated a new server (see below for details) which we decided to make our new science database. Our current science database works just fine, but is on an older (slower) system with a used set of fibre channel disks that frequently fail for one reason or another.
We finally got the new server set up to our liking last week and started unloading all the database tables this week from the old system. The increased disk activity caused the aforementioned disks to completely freak out last night. We had to shut down the science database this morning and we are still in the process of recovering the system.
While most of the BOINC backend functions without any dependency on the science database, the splitters and assimilators do not. The assimilators being off are no big deal - this just means a delay in moving results on disk into the database. But when the splitters are off no new work can be created, and our queue of work to send already ran dry this afternoon.
It is highly unlikely we will get the database back up before the end of the day, or anytime this weekend. Even if we do, our highest priority will be to unload the remaining few database tables before the disks crash again.
If you want to keep your computers busy, you can always work on multiple projects.
- 11.09.2006, 21:37 Seti: Kurze Wartungsarbeiten #14
Update: Eventuell können im Laufe der Nacht (in den Staaten noch Abend) neue WUs generiert werden. Es müssen noch Kleinigkeiten erledigt werden.
September 11, 2006
Outage update: Though we moved a bulk of the data onto the new science database server over the weekend, there is still plenty to do before we can start using this new server. In the meantime, we may be able to start creating new work to send out later today.
- 11.09.2006, 22:45 Seti: Kurze Wartungsarbeiten #15
- 12.09.2006, 12:42 Seti: Kurze Wartungsarbeiten #16September 11, 2006 - 22:00 UTC
We're back on line, more or less, after the science database crash last Friday. Work is being generated, and the assimilators are catching up on the backlogged results. But we're far from out of the woods.
We aren't quite finished getting the data unloaded from the old (broken) database. There is no data corruption, but we are currently operating on only one half of a RAID 10 mirror. In other words, one more drive failure and the whole database is toast. What happened? The two mirrors are on two separate drive arrays, and one of the two enclosures went belly-up, causing all our headaches this past Friday, and our cautionary measures over the weekend.
The rest of the data should be unloaded within the next 24 hours. Then we have to check the data, and load it onto the new server. Then we check to make sure the data on both databases match. And then we shut down the whole project for a day and migrate all the workunits created and results uploaded since we started this whole project a week ago.
Finally, we turn things over to the new database and start the project back up. If all is well, we'll be completely on the new server and can fully retire the old one.
Meanwhile there is positive news about the data recorder. We haven't been able to take much data with the new multi-beam recorder because of DLT drive problems galore. In an effort to move away from that technology, we successfully implemented and tested using swappable SATA drives to store the data. As of this morning we have all the parts working at Arecibo. Soon we will get the parts ordered/installed/tested up here at the lab. Eventually, instead of shipping data back on forth on tapes, we'll be shipping whole 500GB drives.